5_._ _Minimum Efficient Scale and Natural Monopolies

Pandora is a small island with just one power plant, which is a monopolist in the island’s electricity market. The plant’s cost function is ?(?)=0.5?3−20?2 + 282.5 ?, where ? is the number of MW-h of electricity produced per day. This implies the marginal cost curve ?C(?) =1.5 ?2 –40? + 282.50. Daily demand for electricity in Pandora is given by ??(?) = 16 −(1/20)?, where ? is the price of a MW-h of electricity.

a) Find the power plant’s minimum efficient scale. Is this a natural monopoly? Why? [Hint: Can the firm exhaust its economies of scale and still find customers at that output level? You may want to graph the demand curve before you answer that question.]

b) Suppose a private firm built and operates this power plant. In the absence of any government intervention, will the monopolist be maximizing its profit if it charges $220 per MW-h of electricity? Why or why not?

[Hint: Check whether the optimal mark-up formula ([(?−?C)⁄p] = 1/|??|) is satisfied]
Answer #1

a) CC®) = 0.503- 200°+28Q5 Q At minimum efficient scale : MCCQ) = ACCO) av, OC 1502-400 +288.5 = 0.52- 200 +2865 :ar, Q? - 20